OpenStudy (anonymous):

Jillian’s school is selling tickets for a play. The tickets cost $10.50 for adults and $3.75 for students. The ticket sales for opening night totaled $2071.50. The equation 10.50a + 3.75b = 2071.50, where a is the number of adult tickets sold and b is the number of student tickets sold, can be used to find the number of adult and student tickets. If 82 students attended, how may adult tickets were sold?

OpenStudy (anonymous):

Plug in '82' for 'b' in the equation, and solve for 'a': \(10.50a + 3.75b = 2071.50\) \(10.50a + 3.75(82) = 2071.50\) Multiply: \(10.50a + 307.50 = 2071.50\) Sub tract 307.5 to both sides, what's 2071.50 - 307.50? @Rafyy

OpenStudy (anonymous):

1764 lol @iGreen.

OpenStudy (anonymous):

Yes, so we have: \(10.50a = 1764\) Now divide 10.50 to both sides, what's 1764 / 10.50?

OpenStudy (anonymous):

i got 168.

OpenStudy (anonymous):

Yes, so 168 adult tickets were sold.

OpenStudy (anonymous):

\(\bbox [10pt, lime, border:5pt solid black]{\Huge\sf\color{blue}{Welcome\ to\ \color{#00A1FF}{Open}\color{#329932}{Study!!!}}}\) You can give medals by clicking 'Best Response'.